What is the length of the arc on a circle with radius 10 cm intercepted by 20° angle?
Use 3.14 for π
Round the answers to the hundredths place.

which angles are conterminal with 7x/8
-25x/8
23x/8
-9x/8
9x/8

Answer by Alan3354(69443)

What is the length of the arc on a circle with radius 10 cm intercepted by 20° angle?
Use 3.14 for π
----
Convert 20 degs to radians = pi/9
Arc = r*angle = 10pi/9 cm
===============================================

which angles are conterminal with 7x/8

-25x/8
23x/8
-9x/8
9x/8
=========
None of the above.
If it's 7x/8 degs, then it's 7x/8 + n*360 where n = any integer
---
If it's 7x/8 radians, then it's 7x/8 + n*2pi where n = any integer